Given the following market demand and supply function of good X is respectively Qd= 100 - 2P and P = 1/2Qs + 10, find the equilibrium price and equilibrium quantity.


1
Expert's answer
2022-01-11T14:35:59-0500

The equilibrium price and equilibrium quantity are:

Qd = Qs,

Qs = 2P - 20,

100 - 2P = 2P - 20,

P = 30,

Q = 2×30 - 20 = 40 units.